# Explicit World Models for Reliable Human-Robot Collaboration

### v1

@@
Humans learn to interpret the world not only through static visual or speech perception, but through continuous integration of multimodal cues including gaze, gestures, prosody and movement dynamics, and contextual knowledge. These cues carry rich inferential signals about what others mean, what they want, and what will happen next, and contribute to developing a mutual understanding of the world that forms the basis for cooperative action. This shared conception of the world has been termed common ground, a joint understanding of the tasks, communications, and environments between agents (Dillenbourg1999GroundingIM).
@@
The idea of common ground originates from language and cognition studies (Clark1983CommonGA) and has been extensively studied in the field of human-AI teaming under the hood of shared mental models, which cover constructs such as knowledge representation, schema, and situation awareness (Andrews2022TheRO). In the domain of Human-Robot Collaboration (HRC), this concept underpins effective teamwork, requiring sophisticated mechanisms to bridge differences between human and artificial agents in terms of perception, cognition, and embodiment (Tan2020TaskOrientedMQ).
@@
Perceptual grounding is arguably the first step in the establishment of common grounds to construct a valid world model for HRC. Research in this domain has been centred around visual understanding enabled by deep learning models, and more recently visual foundational models. These models have been used to address various tasks and benchmarks on Visual Question Answering (VQA) (Zhong et al. 2022), which nevertheless is inadequate at capturing the dynamic, multimodal, and task-specific context of HRC. Therefore, a growing interest is observed in building common grounds for task-oriented collaborations. We proposed a Task-oriented Collaborative Question Answering (TCQA) benchmark (Tan2020TaskOrientedMQ) for benchmarking grounding methods with quantitative evaluation of their effectiveness in HRC tasks. Our baseline model combining deep learning to tackle basic perception and symbolic reasoning to capture high-level contextual information and reasoning achieved good performance on the benchmark, but this approach still suffers from fragility/errors in novel scenes and lacks flexibility in constructing new semantic inferences. To address these issues, Large Language and Multimodal Models (LLMs/LMMs) have been leveraged for semantic knowledge to inform affordance reasoning (Ahn2022DoAI; Huang2023VoxPoserC3), coordination (Zhang2024TowardsEL) and human goal reasoning (Wan2023HandMeThatHC). However, these approaches face challenges owing to their intrinsic disembodiment from the physical world.
@@
Foundational work in social robotics has emphasised the importance of joint attention and shared intentionality for meaningful interaction. Scassellati demonstrated that joint attention enables robots to interpret human referential cues (scassellati1996mechanisms). Extending this, Breazeal et al. showed that non-verbal behaviours significantly improve efficiency and robustness in human–robot teamwork (breazeal2005effects), revealing that embodied communication is essential for reliable coordination, while Sato et al. showed that continuous monitoring of human behaviours expressed both via conscious actions/language and unconscious/involuntary non-verbal cues is needed for robots to actively infer human intentions (sato1995active). In parallel, work on legible robot motion showed that robots must act not just efficiently, but also expressively, producing behaviours that communicate intent to human partners, improving predictability and coordination in shared workspaces (dragan2013legibility). These works collectively support the argument that reliable collaboration emerges from interactive common ground building, not merely isolated perception. In related work, we demonstrated how multimodal human cues are essential for reliable referential grounding. In M2GESTIC (weerakoon2020gesture), we showed that a distance-weighted understanding of pointing gestures can significantly reduce ambiguity in comprehending natural multi-modal human instructions. We also demonstrated that eye gaze provides strong cues for predicting referents and action steps during joint tasks (johari2021gaze). COSM2IC (weerakoon2022cosm2ic) introduced adaptive real-time multimodal fusion that prioritises gesture or linguistic structure depending on context, highlighting that reliability emerges from dynamic coordination rather than rigid pipelines. Most recently, Ges3ViG (mane2025ges3vig) integrates pointing gestures with 3D visual grounding, advancing spatially grounded reference understanding for real-world embodied AI.
@@
More recently, we observe the emergence of explicit, interpretable, and self-evolving world models as the foundation for next-generation neuro-symbolic intelligence. Weng’s early insights into autonomous mental development (Weng2012AutonomousMentalDev) framed a crucial distinction between symbolic and emergent representations, arguing that genuine intelligence requires the brain—or its artificial counterpart—to develop its internal representations without human handcrafting, forming abstractions directly from sensorimotor experience. In essence, Weng’s vision anticipated the need for agents that construct and continually refine internal world models capable of linking sensory input to motor behaviour and abstract reasoning—a theme that now defines modern neuro-symbolic research.
Consider for example NeSyC (Choi2025NeSyC), a neuro-symbolic continual learner inspired by the hypothetico-deductive model of scientific reasoning. It combines the generative creativity of large language models (LLMs) with the logical precision of symbolic solvers, creating a feedback loop where inductive inference (via LLMs) and deductive validation (via Answer Set Programming) reinforce each other. Through contrastive learning and continual memory refinement, NeSyC can generalise actionable knowledge across diverse open-domain environments — transforming raw experience into structured, symbolic understanding. Our recent work (Nguyen2025KML) on Knowledge Module Learning (KML) and the PKR-QA benchmark for procedural reasoning deepens this trajectory. We encode procedural knowledge in a knowledge graph linking tasks, steps, actions, objects, tools, and purposes, grounding symbolic reasoning in perceptual and temporal context. KML trains neural knowledge modules to capture relations between entities — bridging the gap between statistical learning and symbolic compositionality. When combined with LLM-generated reasoning programs, these modules yield interpretable, stepwise reasoning traces that can be verified and debugged. The integration of structured knowledge with neural embeddings transforms procedural understanding from mere sequence prediction into causal reasoning — enabling agents to explain why a particular action should occur, not just what should be done next.

### v2

@@
Humans learn to interpret the world not only through static visual or speech perception, but through continuous integration of multimodal cues including gaze, gestures, prosody and movement dynamics, and contextual knowledge. These cues carry rich inferential signals about what others mean, what they want, and what will happen next, and contribute to developing a mutual understanding of the world that forms the basis for cooperative action. This shared conception of the world has been termed common ground, a joint understanding of the tasks, communications, and environments between agents (Dillenbourg and Traum 1999).
@@
The idea of common ground originates from language and cognition studies (Clark et al. 1983) and has been extensively studied in the field of human-AI teaming under the hood of shared mental models, which cover constructs such as knowledge representation, schema, and situation awareness (Andrews et al. 2022). In the domain of Human-Robot Collaboration (HRC), this concept underpins effective teamwork, requiring sophisticated mechanisms to bridge differences between human and artificial agents in terms of perception, cognition, and embodiment (Tan et al. 2020).
@@
Perceptual grounding is arguably the first step in the establishment of common grounds to construct a valid world model for HRC. Research in this domain has been centred around visual understanding enabled by deep learning models, and more recently visual foundational models. These models have been used to address various tasks and benchmarks on Visual Question Answering (VQA) (Zhong et al. 2022), which nevertheless is inadequate at capturing the dynamic, multimodal, and task-specific context of HRC. Therefore, a growing interest is observed in building common grounds for task-oriented collaborations. We proposed a Task-oriented Collaborative Question Answering (TCQA) benchmark (Tan et al. 2020) for benchmarking grounding methods with quantitative evaluation of their effectiveness in HRC tasks. Our baseline model combining deep learning to tackle basic perception and symbolic reasoning to capture high-level contextual information and reasoning achieved good performance on the benchmark, but this approach still suffers from fragility/errors in novel scenes and lacks flexibility in constructing new semantic inferences. To address these issues, Large Language and Multimodal Models (LLMs/LMMs) have been leveraged for semantic knowledge to inform affordance reasoning (Ahn et al. 2022; Huang et al. 2023), coordination (Zhang et al. 2024) and human goal reasoning (Wan et al. 2023). However, these approaches face challenges owing to their intrinsic disembodiment from the physical world.
@@
Foundational work in social robotics has emphasised the importance of joint attention and shared intentionality for meaningful interaction. Scassellati demonstrated that joint attention enables robots to interpret human referential cues (Scassellati 1996). Extending this, Breazeal et al. showed that non-verbal behaviours significantly improve efficiency and robustness in human–robot teamwork (Breazeal et al. 2005), revealing that embodied communication is essential for reliable coordination, while Sato et al. showed that continuous monitoring of human behaviours expressed both via conscious actions/language and unconscious/involuntary non-verbal cues is needed for robots to actively infer human intentions (Sato et al. 1995). In parallel, work on legible robot motion showed that robots must act not just efficiently, but also expressively, producing behaviours that communicate intent to human partners, improving predictability and coordination in shared workspaces (Dragan et al. 2013). These works collectively support the argument that reliable collaboration emerges from interactive common ground building, not merely isolated perception. In related work, we demonstrated how multimodal human cues are essential for reliable referential grounding. In M2GESTIC (Weerakoon et al. 2020), we showed that a distance-weighted understanding of pointing gestures can significantly reduce ambiguity in comprehending natural multi-modal human instructions. We also demonstrated that eye gaze provides strong cues for predicting referents and action steps during joint tasks (Johari et al. 2021). COSM2IC (Weerakoon et al. 2022) introduced adaptive real-time multimodal fusion that prioritises gesture or linguistic structure depending on context, highlighting that reliability emerges from dynamic coordination rather than rigid pipelines. Most recently, Ges3ViG (Mane et al. 2025) integrates pointing gestures with 3D visual grounding, advancing spatially grounded reference understanding for real-world embodied AI.
@@
More recently, we observe the emergence of explicit, interpretable, and self-evolving world models as the foundation for next-generation neuro-symbolic intelligence. Weng’s early insights into autonomous mental development (Weng 2012) framed a crucial distinction between symbolic and emergent representations, arguing that genuine intelligence requires the brain—or its artificial counterpart—to develop its internal representations without human handcrafting, forming abstractions directly from sensorimotor experience. In essence, Weng’s vision anticipated the need for agents that construct and continually refine internal world models capable of linking sensory input to motor behaviour and abstract reasoning—a theme that now defines modern neuro-symbolic research.
Consider for example NeSyC (Choi et al. In Press), a neuro-symbolic continual learner inspired by the hypothetico-deductive model of scientific reasoning. It combines the generative creativity of large language models (LLMs) with the logical precision of symbolic solvers, creating a feedback loop where inductive inference (via LLMs) and deductive validation (via Answer Set Programming) reinforce each other. Through contrastive learning and continual memory refinement, NeSyC can generalise actionable knowledge across diverse open-domain environments — transforming raw experience into structured, symbolic understanding. Our recent work (Nguyen et al. 2025) on Knowledge Module Learning (KML) and the PKR-QA benchmark for procedural reasoning deepens this trajectory. We encode procedural knowledge in a knowledge graph linking tasks, steps, actions, objects, tools, and purposes, grounding symbolic reasoning in perceptual and temporal context. KML trains neural knowledge modules to capture relations between entities — bridging the gap between statistical learning and symbolic compositionality. When combined with LLM-generated reasoning programs, these modules yield interpretable, stepwise reasoning traces that can be verified and debugged. The integration of structured knowledge with neural embeddings transforms procedural understanding from mere sequence prediction into causal reasoning — enabling agents to explain why a particular action should occur, not just what should be done next.